What is the maximum vertical height to which a baseball player can throw a ball if he can throw it a maximum distance of 60.0 m? Assume that the ball is released at a height of 1.60 m with the same speed in both cases.

Grade:9

2 Answers

Saurabh Koranglekar
4 years ago
Vikas TU
4 years ago
Dear student 
R = H-h 
R = 60m - 1.60m 
= 58.4 m 
Y max = R/2 
Ymax = 29.2m 
Hmax = ymax + 1.60m 
Hmax = 30.80 m